OpenStudy (errinss):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = ?, b = 12, c = 19 A. 14.03 B. 14.38 C. 14.73 D. 15.07

OpenStudy (errinss):

@satellite73 @rishavraj

OpenStudy (errinss):

@Cheez

OpenStudy (starriebaby):

a

OpenStudy (starriebaby):

is it

OpenStudy (cheez):

A^2 * B^2 = C^2. So, C^2 / B^2 = A^2 Square C and B, divide C by B and you have your answer

OpenStudy (mrnood):

@cheez that is not quite correct in the equation c is the hypotenuse so a^2 + b^2 = c^2 a^2 = c^2 - b^2 you know c and b so use the equation to find a^2, then take square root to find a

OpenStudy (errinss):

i got 2.506

OpenStudy (cheez):

Ah my mistake, I thought it was divided by. I got the 2.5--- answer and was like "ok something wrong...."

OpenStudy (mrnood):

start with c what is c^2 then what is b^2 then what is c^2 - b^2 then take squareroot to get a

OpenStudy (cheez):

It has been like 4 years since I took pre algebra so it takes a toll on ye old memory.

OpenStudy (errinss):

the answer is C